OpenStudy (anonymous):

can someone please help?? what is y=mx + (2+m) in general form??

OpenStudy (saifoo.khan):

y = m(m + x) + 2

myininaya (myininaya):

general form I believe is Ax+By=C y=mx+(2+m) -mx+y=2+m is general form where A=-m, B=1, and C=2+m